OpenStudy (anonymous):

Solve the equation or formula for the indicated variable. S=6r^3t,for t

OpenStudy (anonymous):

\[S=(6*r)^(3*t)\] If you take the natural logarithm of both sides, you get \[\ln(S)=\ln((6*r)^(3*t))\] Take the 3*t to the front of the natural log (we can do this because of properties of logarithms) and we get \[\ln(s)=(3*t)*\ln(6*r)\] Then just divide both sides by \[\ln(6*r)\] and \[\ln(S)\div(\ln(6*r))=3*t\] Divide again by three and \[t=ln(S)\div(3*\ln(6*r))\]

OpenStudy (anonymous):

Thank you:)